Sol Summary – April 25th

Crew 297 Sol Summary Report 25-04-2024

Sol: 11

Summary Title: Earth Rising on the Red Horizon

Author’s name: Pawel Sawicki

Mission Status: Nominal

Sol Activity Summary:

This morning, the first of two EVAs was conducted. Drs. Storch and Sawicki ventured out far north to the Gateway to Lith this morning to locate an area suitable for Case to drive around. Although they did not find anything in the vicinity of Gateway to Lith for this purpose, they admired the views of this area. On the way back, an adequate area was indeed discovered and utilized for the human-research interaction study. Case performed well in its given task, requiring minimal human intervention. The robot was then driven manually to test its steep hill-traversing abilities. The equipment was gathered up and the crew returned home to the Hab.

The second, late afternoon EVA, involved the two Crew Engineers Storch and Laude.This EVA team journeyed to Somerville Overlook to acquire telephoto photos of the sedimentary rock layers in the walls of the escarpments. As the weather started to head south, the team headed back towards MDRS, but made a planned stop at Chandor Chasma for additional photos of stratigraphy and rocks. Terrain was also examined for NPS recovery and was found extremely difficult for such an operation.

For dinner the crew was treated to a flavorful vegetable stir fry prepared by Chef (and GHO) Marquez who used a substantial amount of harvested greens from the GreenHab.

Look Ahead Plan:

Tomorrow is unfortunately our last sol at MDRS. We will start the day with a geological expedition EVA. When we return we will have our hands full with end-of-mission tasks: finishing reports, dismantling the GreenHab for the season, taking inventory, packing for our trip back to Earth and cleaning.

Anomalies in work:

Suit 11 (fan power substantially decreased mid-EVA): Battery was replaced but the air flow is still less than nominal. Inop’d for the current mission.
Morning (data from 0900): calm winds (2 mph), chilly (59F), and scattered cloud coverage
Noon: calm winds (7 mph), warm (70F), and scattered cloud coverage
Afternoon (1800): moderate winds (16 mph), chilly (69F), and scattered cloud coverage
Crew Physical Status: Overall crew health is nominal.

EVA: There were two EVAs today. The first EVA had a successful testing effort of Case on the way back from the scenic Special Region and Gateway to Lith. The second EVA went to Somerville Overlook and Candor Chasma, collecting telephotos of the geological features in those areas for review by the crew geologist.

Reports to be filed:

Sol Summary
Journalist Report
GreenHab Report
Operations Report
EVA report (2 EVAs)
EVA request (1 EVA)
Support Requested: None requested.

Copyright © The Mars Society. All rights reserved. | Main Site